Teach me the steps of solving a ratio problem such as 2:3:4


Answer by Theo(13342)

here's a good reference.

http://www.onlinemathlearning.com/ratio-math-problems.html

just do a search for "how to solve ratio problem with 3 ratios" and you'll find lots more websites that will attempt to explain it.

to use your question as an example:

assign variables to the ratios

i'll use one of the examples in the reference i chose to show you how it works.

Clothing store A sells T-shirts in only three colors: red, blue and green. The colors are in the ratio of 3 to 4 to 5. If the store has 20 blue T-shirts, how many T-shirts does it have altogether?

the colors are red, blue, and green.

the colors are in the ratio of 3 to 4 to 5.

this means that the ratio of red shirt to blue shirts is 3/4 and the ratio of blue shirts to green shirts is 4/5

you get:

red/blue = 3/4 and you get blue/green = 4/5

you have 20 blue shirts.

your ratio of red/blue = 3/4 = x/20

x represents the number of red shirts you need to find.

solve for x to get x = 20*3/4 = 60/4 = 15

your ratio of blue/green = 4/5 = 20/x

x represents the number of green shirts you need to find.

solve for x to get x = 5*20/4 = 25

you have 15 red shirts, 20 blue shirts, and 25 green shirts.

15/20 = 3/4

20/25 = 4/5

if you simplify 15/20, you will see that it is equivalent to 3/4

if you simplify 20/25, you will see that it is equivalent to 4/5.

the total number of shirts is 60.

check the reference and you'll see that this is correct.

here's another tutorial that might be helpful.

http://www.purplemath.com/modules/ratio2.htm